The principal value of the argument of

(1+i√3)(1−i) is____



1
Expert's answer
2021-05-24T03:16:02-0400
"(1+3i)(1-i)=1-i+3i+3=4+2i"

"x=4>0, y=2"

"\text{Arg}(4+2i)=\tan^{-1}\big(\dfrac{2}{4}\big)"



"\text{Arg}((1+3i)(1-i))" is "\tan^{-1}\big(\dfrac{1}{2}\big)."